Kyt:

most versoza wants to compare simple and compound interest on a 3,500 pesos investment for three years find the interest if one's current 6.5% simple interest annually find the interest if found in 6.5% interest compounded annually find the difference between the two interest

itsme:

whut r ur thoughts on this 1 ? whut do yu think we hav to do first?

itsme:

@Kyt

Mercury:

some variables to define: A = final amount P = principal (the starting amount) r = interest rate as a decimal t = time (in years) simple annual interest: A = P(1+rt) compound annual interest: A = P(1+r)^t plug in the appropriate quantities, evaluate both final balances, then subtract (larger balance) - (smaller balance)